Red Beads cost $1.30 per box, they are mixed with yellow beads that cost $2.00 a box. How many boxes of each color must be used to make 10 boxes of beads costing $1.72 per box?
Answer by DeepaEvangeline(15)

Cost of red bead box = $1.30
Cost of yellow bead box = $2.00
Let number of red beads boxes be x
Let number of yellow beads boxes be y
Total number of boxes = 10
Cost of each box = $1.72
Cost of 10 boxes = 1.72 x 10 = $17.2
The equations are
x + y = 10 -----------------(1)
1.3x + 2y = 17.2 -----------(2)
Solving the equations (1) and (2) gives
Multiply eqn (1) by 2
2x + 2y = 20 ---------------(3)
Eqn (3) - (1) gives
2x - 1.3x + 2y - 2y = 20 - 17.2
.7x = 2.8
x = 2.8/.7 = 4
Therefore x + y = 10
4 + y = 10
y = 10 - 4 = 6
Number of red beads boxes is 4
and number of yellow beads boxes is 6